here's the question- The sum of 20 and a number is 30?
Answer by elima(1433)

The sum of 20 and a number is 30?
The sum - means it is an addition problem, a number we will say is x, is means =
20 + x = 30
subtract 20 from both sides;
20-20+x=30-20
x=10